OpenStudy (anonymous):

X,y,m,n are real numbers that: 2^x.3^y=2^m.3^n Prove x=m and y=n

OpenStudy (loser66):

i don't know !! it's trivial

OpenStudy (anonymous):

Can you help me prove this execrise.

OpenStudy (firekat97):

is this the question ? \[2^x. 3^y = 2^m. 3^n\]

OpenStudy (anonymous):

Not sure if this is a trick question or not, but by using properties of index laws. We can say that if a particular "base" and it's exponential is equal to some unknown variable. I think you can say that in this example, that: 2^x.3^y = 2^m.3^n x = m and y = n Provided that you say, by Index laws...

OpenStudy (anonymous):

@FireKat97: yes